OpenStudy (amtran_bus):

Help me solve 4x-x^2 =0

hartnn (hartnn):

factor out x, what u get ?

OpenStudy (amtran_bus):

\[-(x(x+4))=0\]

hartnn (hartnn):

and since you factored out - also, the +4 becomes -4

OpenStudy (amtran_bus):

Ohh, I see. Was there another way to factor the -?

hartnn (hartnn):

x(4-x) = -x (x-4)

hartnn (hartnn):

nopes, this is correct

OpenStudy (amtran_bus):

So x=0 and x=4. Thanks.

hartnn (hartnn):

correct :) welcome ^_^